OpenStudy (medina13):

Which expression is equal to Which expression is equal to. X 4/3 * x 1/3 ? Answers. X 2/3 X 5/3 X 5/9 X 4/9

jimthompson5910 (jim_thompson5910):

Rule: \[\Large x^a*x^b = x^{a+b}\] Example of this rule in action \[\Large x^{7/9}*x^{5/9} = x^{7/9+5/9}\] \[\Large x^{7/9}*x^{5/9} = x^{12/9}\] Recall that to add fractions the denominators must be the same and we add the numerators over the common denominator.